a body falls freely from rest covers more distances in each next second .these distances are in ratio?

Grade:12th pass

2 Answers

Khimraj
7 years ago
Distance traveled in each nth sec is given bySn = (1/2)g(2n-1)So the ratio of distance in 1st sec 2nd sec 3 sec and so on is given by1:3:5:7:9...
Nikhil Sukhani
7 years ago
Let u = initial speed 
      t = time taken
      a = acc. 
then ,
 distance covered in n seconds (S1) = u(n) +1/2a(n)2
                                                                     = nu + 1/2an2
 distance covered in n-1 seconds (S2) = u(n-1) +1/2a(n-1)2
                                                                           = nu - u + 1/2an+1/2a - an 
  distance vovered in nth second = Sn – Sn-1
                                                              = nu +1/2an- ( nu - u +1/2an2 +1/2a – an )
                                               = u – 1/2a +an
                                               = u -a(1/2-n)
                                               = u + a/2(2n-1)
Since body falls freely from rest
Therefore ,  a=g
                  u=0
So , the formula becomes g/2(2n-1) 
So the ratio of distance travelled in 1st sec , 2nd sec 3rd sec.... is 1:3:5:...
CHEERS ;)